OpenStudy (godlovesme):

help pweaseee D: Andy has two similar conical hats, Hat A and Hat B. The diameter of Hat A is 18 inches, and the diameter of Hat B is 9 inches. What is the ratio of the volume of Hat A to the volume of Hat B? 2:1 16:1 4:1 8:1

OpenStudy (nuttyliaczar):

Volume of a cone is \[\pi*r ^{2}*h/3\]

OpenStudy (nuttyliaczar):

And diameter is just 2r, so for hat A r=9 in and hat B r=4.5 in

OpenStudy (godlovesme):

im following :)

OpenStudy (nuttyliaczar):

Set up an equation that has volume on both sides like so: \[\pi*r _{A}^{2}*h/3=\pi*r _{B}^{2}*h/3\]

OpenStudy (nuttyliaczar):

As you can see, the pi and h/3 will cancel out once you divide both sides by them. Then you are left with r^2 on each side for the respective cones.

OpenStudy (nuttyliaczar):

And you can take it from there =P

OpenStudy (godlovesme):

ok one sec :D

OpenStudy (godlovesme):

so it would be.. \[9^2 = 4.5^2\] which will simplify to \[4:1\] ? :D

OpenStudy (anonymous):

smarty pants :3

OpenStudy (godlovesme):

X'D

OpenStudy (godlovesme):

thanks guys :))

OpenStudy (nuttyliaczar):

Yeah you got it!

OpenStudy (nuttyliaczar):

Every ratio type problem can be handled this way

OpenStudy (godlovesme):

alright thank you so much :))

OpenStudy (godlovesme):

it wasnt 4:1 :( .. oh well thanks though :)

OpenStudy (godlovesme):

XDD

OpenStudy (anonymous):

:O oha! how so?

OpenStudy (nuttyliaczar):

Hmm is it not? Let me plug everything in myself

OpenStudy (godlovesme):

no idea @annas and @nuttyliaczar no worries its fine , unless u want to check it :3

OpenStudy (nuttyliaczar):

I just checked and still got 4:1

OpenStudy (godlovesme):

i'll asj my teacher, thanks anyways :)

OpenStudy (godlovesme):

annas cookies *_*

OpenStudy (nuttyliaczar):

Volume for A = pi*r^2*h/3 = 254.469*h/3 Volume for B = pi*r^2*h/3 = 63.617*h/3

OpenStudy (nuttyliaczar):

Dividing them gives 4/1 so... I'm not sure why that was wrong

OpenStudy (godlovesme):

its ok i'll talk to my teacher :) thanks though i really appreciate ur help guys :)